OpenStudy (anonymous):

Write the ratio in three different formats: to , , and . Carlos biked 24 miles and ran 6 miles. Write the ratio of the distance Carlos biked to the total distance.

OpenStudy (anonymous):

to , , and . ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

"Write the ratio in three different formats: to , , and . "

OpenStudy (anonymous):

Just any three formats?

OpenStudy (anonymous):

can you tell me one

OpenStudy (anonymous):

Well we can write it as a proportion such as\[\frac{ 2 }{ 1 }\]

OpenStudy (anonymous):

A fraction, I don't know why I said a proportional

OpenStudy (anonymous):

is it 6 to 25 or a3 to 12 or 30 to 24 or 4 to 5

OpenStudy (anonymous):

i skink at math

OpenStudy (anonymous):

Are these the possible answers or ideas?

OpenStudy (anonymous):

possible answers

OpenStudy (anonymous):

Okay :) So we have the ratio: \[\frac{ 24 }{ 6 }\]

OpenStudy (anonymous):

But we should be able to simplify this

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

They might want the simplified answer though

OpenStudy (anonymous):

to what

OpenStudy (anonymous):

The ratio \[24:6\]

OpenStudy (anonymous):

\[\frac{ 6 }{ 24 }\]

OpenStudy (anonymous):

What's the greatest common factor of 24 and 6?

OpenStudy (anonymous):

The largest number that 24 and 6 can be divided evenly by.

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6

OpenStudy (anonymous):

one more please

OpenStudy (anonymous):

How much grapefruit juice is needed to make 32 servings of fruit punch?

OpenStudy (anonymous):

How much grapefruit juice is needed to make 32 servings of fruit punch?

OpenStudy (anonymous):

For the first question, since the ratio can be divided by the greatest common factor of 6, the simplified answer would then be \[\frac{ 24 \div 6 }{ 6 \div 6 } = \frac{ ? }{ ? }\]